OpenStudy (anonymous):

Write an equation of the line that pases through the points (6,-1) and (8,3)

OpenStudy (anonymous):

You can use here two point formula..

OpenStudy (anonymous):

\[\large \frac{y- y_1}{y_2 - y_1} = \frac{x - x_1}{x_2 - x_1}\]

OpenStudy (anonymous):